Find the length of the curve, 2y² = x³ from the vertex (0,0) to the point (4,4√2).
1
Expert's answer
2019-10-23T14:54:06-0400

Solution:

We are going to find the length of the curve.



Given curve is

"2 y^2 = x^3"

We can write this for y;

"y = ± \frac {1} {\sqrt 2} x^{\frac {3}{2}}"

Differentiate with respect to x,



"\frac {dy}{dx} = ± \frac {1} {\sqrt 2}(x^{\frac {3}{2}})^{\prime}"

"= ± \frac {1} {\sqrt 2} \times \frac {3} {2} \times x^{\frac {1}{2}} == ± \frac {3} {2\sqrt 2} \times x^{\frac {1}{2}}"

We can find the length of the given curve "2 y^2 = x^3" from (0, 0) to the point (4,4√2) by using a formula.

That is,


Length of the curve =

"=\int_0^ 4 \sqrt {(1+ (\frac {dy}{dx})^2 } \space dx"




"= \int_0^ 4 \sqrt {(1+ (± \frac {3} {2\sqrt 2} \times x^{\frac {1}{2}})^2 } \space dx"



"= \int_0^ 4 \sqrt {(1+ ( \frac {9} {8} \times x) } \space dx"

"= \frac {1} {2\sqrt 2 } \int_0^4 \sqrt {8+ {9} x } \space dx"

"Let \space 8 +9x = t\\ 
 9 dx = dt \\
 dx = \frac {1}{9} dt"

"Limits : x = 0 \space then \space t = 8 \\
 x = 4 \space then \space t = 44"

"= \frac {1} {2\sqrt 2 } \int_8^{44} \sqrt t \space \frac {1}{9} \space dt"

"= \frac {1} {18\sqrt 2 } \int_8^{44} t^{\frac {1}{2}} \space \space dt"

"= \frac {1} {18\sqrt 2 } \space [\frac {t^{\frac {1}{2} + 1}} {\frac {1}{2} +1}]_8^{44}"

"= \frac {1} {18\sqrt 2 } \space \frac {2}{3} ((44)^{\frac {3}{2}} - (8)^\frac{3}{2})"

"= \frac {1} {27\sqrt 2 } (88\sqrt {11} - 16 \sqrt 2)"

Answer:

"Length \space of \space the \space curve= \frac {1} {27\sqrt 2 } (88\sqrt {11} - 16 \sqrt 2)"